Circular No. 8547

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2005 M1 (CHRISTENSEN) E. J. Christensen, Lunar and Planetary Laboratory, reports his discov-

ery of a comet on images taken with the 1.5-m f/2 reflector in the course of the Mount Lemmon Survey (discovery observation given below). On four co-added 90-s CCD images from June 17.4 UT, Christensen noted a 5′′ coma and a 10′′ tail in p.a. 230^{o}–240^{o}; images taken in improved seeing on June 18.4 showed a 15′′ tail in p.a. 230^{o} and a compact coma of diameter ≈ 7′′ (four co-added 60-s exposures). Following posting on the ‘NEO Confirmation Page’, A. C. Gilmore found a moderately condensed, circular coma of diameter ∼ 15′′ (but no tail) on four co-added 120-s CCD frames taken on June 19.7 with the 1.0-m f/7.7 reflector at Mount John University Observatory.

2005 UT  α_{2000}  δ_{2000}  Mag.
June 17.41206  22^{h}41^{m}08.^{s}59  −7^{o}52′07′′.8  19.9

The available astrometry, the following very preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2005-M33.

T = 2005 Jan. 24.728 TT  ω = 140.^{o}756
Ω = 143.962 }2000.0
q = 2.94607 AU  i = 8.741

COMETS C/2005 G7, C/2005 H2, AND C/2005 H3 (SOHO) Following are the initial positions for additional comets found on SOHO

website images (continuation to IAUC 8546); C/2005 H2 belongs to the Meyer group, the others being Kreutz sungrazers. K. Battams writes that C/2005 G7 had a diffuse ‘teardrop’ shape, reaching mag ∼ 7.5 at 7.1R_{⊙} (Apr. 14.085 UT). C/2005 H2 was elongated and faint, but Battams notes that this may or may not have been a tail, adding that it is not uncommon to see stars appear “diagonally elongated” in the SOHO images; it peaked in brightness at mag ∼ 7.8 at 7.0R_{⊙} (Apr. 17.726). C/2005 H3 was partly obscured by the pylon in the C3 coronagraph, but in C2 images it reached mag ∼ 5.2 at 6.6R_{⊙} on Apr. 22.368, when its tail length was 93′′.

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 G7 Apr. 14.171 1^{h}38.^{m}4 + 8^{o}36′ C2  HS 2005-K43
C/2005 H2 17.629 1 48.2 +12 06 C2 HS 2005-K43
C/2005 H3  21.779 2 14.4 +10 16 C3/2 HS 2005-K43
2005 June 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green